### v2

# Hits to Higgs: Reconstruction-Free Higgs Classification from Raw LHC Detector Data Using Higgsformers

Abstract
We present a comparative study of Higgs event classification at the Large Hadron Collider that bypasses the traditional reconstruction chain. As a benchmark, we focus on distinguishing $t\bar{t}H$ from $t\bar{t}$ events with $H\to b\bar{b}$, a particularly challenging task due to their similar final-state topologies. Our pipeline begins with event generation in Pythia8, fast simulation with ACTS/Fatras, and classification directly from raw detector hits.
We show for the first time that a transformer model originally developed for inner tracker hit-to-track assignment can be retrained to classify Higgs events directly from raw hits. For comparison, we reconstruct the same events with Delphes and train object-based classifiers, including multilayer perceptrons and the Particle Transformer. We evaluate both approaches under varying dataset sizes and pileup levels. Although Higgsformer works exclusively with inner tracker hits (i.e., without calorimeter or muon information), it achieves strong performance with an AUC value of 0.792.
1 Introduction
The Large Hadron Collider (LHC) at CERN produces proton–proton collisions at an unprecedented scale, up to 40 million events per second, generating hundreds of terabytes of online data and tens of petabytes of stored offline data annually. Traditionally, this massive volume is processed through deeply structured reconstruction pipelines. These include particle trajectory fitting, calorimeter clustering, and jet and object identification, which ultimately reduce the raw detector output into a compact set of high-level observables used in physics analyses. While effective, this approach imposes strong inductive biases and may discard low-level information that could be critical for certain tasks.
This raises a natural question: Can modern machine learning models learn directly from the raw detector data, bypassing intermediate reconstruction stages and high-level features altogether? In this work, we address this question by directly comparing two pipelines: one based on an end-to-end classifier operating on raw digitized detector hits obtained with ACTS/Fatras, and another using a baseline classifier trained on high-level reconstructed objects obtained from a Delphes-simulated detector. Both pipelines share identical physics event generation to ensure a fair comparison.
The classification task we consider is the separation of $t\bar{t}$ from $t\bar{t}H$ events, with the Higgs boson forced to decay to a $b\bar{b}$ pair. Discriminating $t\bar{t}H$ events from the dominant $t\bar{t}$ background, especially in the $H\to b\bar{b}$ decay channel, is a challenging tasks in Higgs physics at the LHC ATLAS:2024gth ; CMS:2024fdo .
This process pair provides a realistic and challenging benchmark: they share similar final-state topologies, but differ subtly in object multiplicities and kinematics, particularly in the presence of additional $b$-jets from the Higgs decay. These subtle differences make it an ideal case study for assessing whether low-level hit information can rival or surpass traditional high-level approaches.
We build a complete end-to-end event simulation pipeline based on Pythia8 Sj_strand_2015 event generation, ACTS/Fatras Spannagel:2019actf detector simulation, and a FlashAttention-accelerated dao2023flashattention2 transformer model to classify events directly from hit patterns. A closely related model was previously developed by us as a Trackformer, used for assigning hits to tracks  caron2025 . In this work, we adapt the same basic architecture for Higgs event classification and refer to it as the Higgsformer. It is a lightweight, set-based Transformer designed to operate directly on raw detector hits.
For comparison, we construct a parallel baseline where the same generated events are processed through Delphes deFavereau:2013fsa to reconstruct standard physics objects (jets, leptons, photons, and missing transverse energy). On this object-level representation, we benchmark against multilayer perceptron (MLP) baselines and the Particle Transformer (ParT) Qu:2022mxj . This setup allows us to highlight the novelty of hit-level end-to-end learning with the Higgsformer, while still enabling a comparison to established object-level approaches. However, this comparison should be interpreted with care, since in this prototype study we contrast a full detector setup, which includes the calorimeter, muon system, and inner tracker, with models that operate only on inner tracker data.
We then evaluate both pipelines under varying detector conditions, including datasets with and without additional proton–proton interactions (pileup), to assess robustness in realistic LHC-like environments.
We find that the hit-level model achieves performance competitive with, and in some cases approaching, that of the reconstructed-object baseline, despite bypassing the entire reconstruction chain. This result supports the potential of hit-level end-to-end learning in high-energy physics.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 describes the event generation procedure and the two parallel datasets: the hit-level dataset obtained with ACTS/Fatras and the reconstructed-object dataset obtained with Delphes. Section 4 introduces the machine learning architectures employed in this study, with particular emphasis on our novel Higgsformer model. Section 5 presents the experimental setup and comparative results, covering performance as a function of dataset size, pileup conditions, and learned feature exploration. Finally, Section 6 summarizes our conclusions andand outlines directions for future work.
2 Related Work
The classification of $t\bar{t}H$ events particularly in the $H\rightarrow b\bar{b}$ decay channel has been the focus of various machine learning (ML) efforts in high-energy physics. Prior work has explored both traditional algorithms and neural approaches, typically relying on handcrafted features derived from reconstructed-level data Aaboud_2018 ; Sirunyan_2018 .
Våge vage2018 examined the performance of Boosted Decision Trees (BDTs) and Neural Networks (NNs) in the context of $t\bar{t}H(b\bar{b})$ classification at ATLAS. Using standard high-level features such as jet multiplicities, b-tag scores, and invariant masses, the study found comparable results between BDTs and fully connected NNs, both achieving area under the ROC curve (AUC) values in the range of 75–77%. However, it also highlighted that neural networks are more sensitive to class imbalance and training biases, especially in low-statistics regions.
Similarly, Santos et al. santos2017 conducted a broad benchmarking study of several ML classifiers, ranging from K-Nearest Neighbors and Naive Bayes to XGBoost and neural network variants, on simulated $t\bar{t}H$ events with $H\rightarrow b\bar{b}$. Their best models (XGBoost and NeuroBGD) achieved AUCs around 80.0–80.2% and F-scores around 74%, but all required pre-engineered input variables and physics-specific observables.
While these studies demonstrated the power of ML in this classification task, they all operated on reconstructed objects such as jets, leptons, and missing transverse energy (MET). That paradigm implicitly assumes the availability of full event reconstruction, and carries potential biases from detector modeling, jet clustering algorithms, or b-tagging procedures.
In contrast, our work takes a radically different approach. We eliminate the reconstruction step entirely and operate directly on the raw detector hits using transformer-based models. Inspired by recent work on TrackML caron2025 , we hypothesize that deep architectures can learn to extract task-relevant structure directly from hit patterns given enough training data and realistic detector simulation. To the best of our knowledge, this is the first study exploring end-to-end classification of $t\bar{t}$ vs. $t\bar{t}H$ events from hit-level data alone, without using any handcrafted features or reconstructed quantities.
3 Data Preparation
3.1 Event Generation
We simulate signal and background processes using Pythia8 integrated within the ACTS framework.
For the signal, we consider $t\bar{t}H$ events where the Higgs boson is forced to decay into a $b\bar{b}$ pair; a dominant decay channel for $m_{H}\approx 125$ GeV and of high relevance for experimental analyses. We include both gluon-gluon fusion and quark-antiquark annihilation production channels:
$gg\rightarrow t\bar{t}H,\quad q\bar{q}\rightarrow t\bar{t}H\,.$
Initial and final state radiation (ISR/FSR) are enabled to simulate realistic parton showering. We apply a parton-level transverse momentum cut of $\hat{p}_{T}>100$ GeV to suppress soft events. Hadron-level decays are fully enabled for both the Higgs and top quarks.
The background consists of standard $t\bar{t}$ production, including QCD-induced additional jet radiation. This class includes light-flavor jets but lacks the explicit $b\bar{b}$ pair from Higgs decay.
Each event is generated with Gaussian vertex smearing to reflect LHC beam spot properties:
$\sigma_{x,y}=12.5~\mu\mathrm{m},\quad\sigma_{z}=55.5~\mathrm{mm},\quad\sigma_{t}=5~\mathrm{ns}$
A fixed random seed of 42 is used for reproducibility. All datasets are produced in both ROOT and CSV formats, with 20,000 signal and 20,000 background events per configuration.
3.2 Pileup, postprcessing and reconstruction pipeline
We use the Fatras fast simulation engine provided by ACTS to simulate the interaction of particles with a tracking detector under a 2 T magnetic field. Particle propagation through the detector geometry includes effects such as multiple scattering, energy loss, and material interaction. The detector geometry is defined using the ACTS GenericDetector, mirroring the configuration used in the TrackML challenge kiehn2019trackml . Fatras outputs truth-level hits (SimHits) for each detector surface where a particle intersects. Following simulation, we perform digitization to model sensor response. Digitization is configured using ACTS JSON-based configuration files to emulate realistic pixel and strip detector behavior. The result is a set of digitized hits stored in CSV format, suitable for downstream processing and machine learning input.
To create TrackML-like datasets, we first perform a coordinate transformation from local sensor coordinates to global $(x,y,z)$, then tag each hit with its volume, layer, and module indices. We also compute the transverse momentum, count the number of hits per track ($n_{\text{hits}}$), and remove zero-charge particles or malformed hits. Finally, inspired by TrackML, we assign per-hit weights in the same manner.
We generated datasets with pileup (PU) levels of 0, 5 and 20 interactions per bunch crossing, each with 40,000 events.  [^1] The full dataset pipeline is available at https://github.com/EugeneShalli/hits-gen. This enables evaluation under varying detector occupancy and complexity.
For the baseline comparison, we generate a parallel dataset using the same Pythia8 configuration described in Section 3.1, ensuring identical signal and background definitions, decay channels, and parton-level cuts. This guarantees a fair comparison between the hit-level and object-level approaches. Detector effects are simulated using Delphes 3.5.0 deFavereau:2013fsa with the default ATLAS detector card. The simulation models the calorimeter, tracking, and muon systems, including realistic resolutions, efficiencies, and acceptance cuts. Jets are reconstructed using the anti-$k_{T}$ algorithm with $R=0.4$, and standard ATLAS-like $b$-tagging efficiencies and mis-tag rates are applied. The output consists of reconstructed high-level physics objects:
  • Electrons and muons with four-momenta and isolation variables.
  • Jets with $p_{T}$, $\eta$, $\phi$, energy, and $b$-tag discriminants.
  • Missing transverse energy and its azimuthal angle.
Objects are ordered by decreasing $p_{T}$ and padded to a fixed multiplicity per category to enable batch processing in the machine learning pipeline. Missing entries are zero-padded, and a corresponding mask array is stored for each event.
4 Machine Learning Models
We introduce Higgsformer, a lightweight set-based Transformer operating directly on raw hit coordinates $(x,y,z)$. Each hit is treated as a token, embedded into a $d$-dimensional space, and processed with $L$ Transformer encoder layers. Masked mean pooling produces an event-level representation, followed by a linear classifier for binary discrimination.
By default, we use the small configuration: $L=2$ layers, $H=2$ attention heads, $d=16$ hidden size. Larger settings up to $L=4$, $H=4$, $d=64$ were also tested. We employ exact FlashAttention dao2022flashattention for memory-efficient self-attention and mixed precision training (AMP) for speed. Dropout (0.3) and weight decay are applied for regularization.
Training uses binary cross-entropy with logits. We found out that logits are background biased despite the balanced dataset, so we incorporate a positive-class weight of 1.5 to balance classes. We use AdamW optimizer, learning rate $10^{-3}$, batch size 64, and early stopping with patience of 100 epochs. Models are trained up to 500 epochs.
Multilayer perceptron (MLP)
As a baseline, we train an MLP Rumelhart:1986gxv on fixed-length feature vectors assembled from Delphes objects. Up to four jets ($j$), four $b$-jets ($b$), and one particle of each of the remaining types (electron, positron, muon, and anti-muon) are selected in descending $p_{T}$ order. Their $(E,p_{T},\eta,\phi)$ values are flattened into a one-dimensional array. We consider two input configurations: (1) all reconstructed objects together with the MET and its azimuthal angle, which effectively captures the response of the full detector; and (2) only $b$-jets, which provides information primarily associated with the inner tracker, while excluding light jets, leptons, photons, and MET.
Features are normalized using statistics from the training set: $E$, $p_{T}$, and MET via Z-score, and $\eta$, $\phi$ via Min–Max scaling. The MLP has three hidden layers (256-128-64 units, ReLU activations) and a sigmoid output. It is trained using Adam optimizer and binary cross-entropy loss, with early stopping based on validation AUC.
Particle Transformer (ParT)
To benchmark against a state-of-the-art object-level model, we employ the Particle Transformer Qu:2022mxj for the $t\bar{t}H$ vs. $t\bar{t}$ classification task. The model operates on a sequence of reconstructed objects, using self-attention layers to capture correlations between particles. The detailed network configuration and hyperparameters follow our previous work Builtjes:2022usj . In this study, we use a simplified version of ParT without interaction matrices or coupling terms.
5 Experiments and Results
Performance at Dataset Sizes (Pileup 0)
We trained Higgsformer-small on PU 0 datasets of 10k, 20k, and 40k events with identical validation/test splits (1k/1k). We evaluate Higgsformer againts the Baseline classifier which uses ‘number of hits per event’ as a seperating feature only. Test ROC AUC increases monotonically with training size for Higgsformer (Fig. 1), consistent with scaling trends observed in Trackformers caron2025 .
Figure 1: ROC AUC curves on Test set for models trained on datasets with increasing size (10k, 20k, 40k).
Performance at Pileup Level (Pileup 0, 5, and 20)
We train Higgsformer on the 38k dataset with the varying pileup of 0, 5 and 20. Performance degrades with pileup yet remains well above random; logit histograms (Fig. 2) show that the model still learns to separate classes even at PU 20, while an $n_{\text{hits}}$-only baseline weakens dramatically with increasing pileup (Fig. 3).
(a) PU 0 (AUC =0.779=0.779)
(b) PU 5 (AUC =0.731=0.731)
(c) PU 20 (AUC =0.654=0.654)
Figure 2: Classification output (Logit) histograms for pileup levels 0, 5 and 20. Signal $t\overline{t}h$ is shown in red and background $t\overline{t}$ in blue.
(a) PU 0 (AUC =0.621=0.621)
(b) PU 5 (AUC =0.560=0.560)
(c) PU 20 (AUC =0.530=0.530)
Figure 3: ROC curves for a counts-only classifier using $n_{\text{hits}}$ at different pileup levels for all 40k events.
Geometric Data Augmentation
We exploit detector symmetries via online $(x,y)$ $\phi$-rotations and a $z\!\to\!-z$ flip (train split only; val/test unchanged).
Table 1: Effect of augmentation on Test ROC AUC for Higgsformer-small 38k, pileup=0. +Aug shows the improvement due to data augmentation
Dataset size Higgsformer-small + Aug $\boldsymbol{\Delta}$
10k 0.704 0.721 +0.017
20k 0.757 0.764 +0.007
40k 0.779 0.792 +0.013
Augmentations lead to consistent gains, aligning the training distribution with known cylindrical and forward–backward symmetries as shown in Table 1.
Comparison to Reconstruction-Object Baselines
We compare our hit-level model to several classifiers trained on reconstructed objects. These include two MLPs with varying inputs trained on fixed-length arrays, and a ParT model that operates on variable-length sequence of physics objects derived from Delphes. As shown in Table 2 and Fig. 4, object-based models achieve high AUC values and show signs of saturation with increasing data, while the hit-level model Higgsformer-small continues to improve with scale. Each model is trained on a dataset of fixed size (10k, 20k, or 40k events), balanced between signal and background. From each dataset, $1\,\mathrm{k}$ events are reserved for validation and $1\,\mathrm{k}$ for testing; the remainder is used for training, denoted as $n_{\text{train}}$. The table reports results at three training sizes, denoted $n_{\text{train}}$ (in thousands): 8, 18, and 38.
Table 2: Test AUC as a function of training set size for object-based modules (MLP, ParT) and the hit-based model (Higgsformer-small). Object-based modules are trained on reconstructed physics objects, while the Higgsformer-small operates directly on raw detector hits.
Model Input AUC @ $n_{\text{train}}$ (k)
8 18 38
MLPa (full detector) Reco objects 0.960 0.951 0.959
MLPb (only $b$-jets) 0.859 0.834 0.848
ParTc (full detector) Reco objects 0.972
Higgsformer-small (only inner tracker) Raw detector hits 0.704 0.757 0.779
Higgsformer-small (augmented, only inner tracker) 0.721 0.764 0.792
[…]
b Includes only 4 $b$-jets; no jets, no leptons, no photons, or MET.
c Includes the full reconstructed particle set (jets, $b$-jets, leptons, photons, and MET).
Figure 4: AUC as a function of training set size for different model architectures. Higgsformer-small shows continues improvement with scale. The horizontal dashed line indicates the baseline performance at $\mathrm{AUC}=0.616$.
Learned Feature Exploration
A counts-only classifier based on $n_{\text{hits}}$ provides limited discrimination that diminishes with pileup, while Higgsformer maintains a substantial advantage (Figs. 2 and 3, Table 3).
Table 3: AUC of $n_{\text{hits}}$-only baseline vs. Higgsformer-small.
Pileup AUC ($n_{\text{hits}}$ only) AUC (Higgsformer) Gain
0 0.621 0.765 +0.144
5 0.560 0.731 +0.171
20 0.530 0.654 +0.124
Crucially, we verify that the model learns Higgs-related structure on 98 unseen events. Using a leave-one-hit-out importance measure and a HepMC → ROOT → CSV matching chain to tag hits from final Higgs descendants, we find that Higgs-tagged hits carry systematically higher importance than non-Higgs hits, with the effect strengthening at larger training size which is shown in Table 4.
Table 4: Higgs and non-Higgs hits importance with Higgsformer-small 38k vs. 8k augmented data.
Metric 38k augmented 8k augmented
Mean avg(H) 0.005777 0.002217
Mean avg(non-H) 0.004560 0.001965
[…]
Events with avg(H) >> avg(non-H) 72/98 51/98
The 3D distributions of the ten most important hits per event (Fig. 5), ranked by their contribution to the final loss, are analyzed to examine how the model captures the symmetries of the system. A rotational symmetry in the x–y plane is expected. As shown in Fig. 5, Higgsformer progressively learns these detector symmetries as the size of the training dataset increases.
(a) Top-1010 important hits for Higgsformer-small 10k.
(b) Top-1010 important hits for Higgsformer-small 40k.
Figure 5: Top-1010 important hits (3D) for Higgsformer-small trained with 10k (left) and 40k (right) training data.
Overall, these results indicate that Higgsformer exploits physically meaningful geometric information and increasingly focuses on hits originating from the Higgs decay as the amount of training data increases.
Time Complexity
Finally, we evaluate the runtime performance of the models. While classical CPU-based tracking typically requires about 1 s per event, Higgsformer-small achieves inference times below 10 ms on an NVIDIA H100 GPU for pileup = 0; representing several orders of magnitude speed-up.
6 Conclusions
Our results demonstrate that the Higgsformer model, trained solely on inner tracker hits, can effectively separate $t\bar{t}H$ from $t\bar{t}$ events, achieving a test ROC AUC of 0.792 at zero pileup. Performance is likely limited by the size of the training dataset and the degree of model optimization.
The analysis of learned features in Section 5 reveals that the network does not rely on simple proxies such as the total number of hits. Instead, it systematically assigns higher importance to hits traceable to Higgs decay products. With increasing training data, high-importance hits form more symmetric and cylindrical patterns, suggesting that Higgsformer learns both detector geometry and Higgs-specific structures.
While object-level models (based on the full detector) still outperform Higgsformer (which uses only inner tracker inputs) in this study, our results show that modern attention-based architectures can extract meaningful event-discriminative information directly from raw hit-level data. This represents a promising step toward reconstruction-free classification in high energy physics.
In addition to improving the ML model architecture, future work will focus on scaling to larger datasets, integrating additional subdetectors, and testing the approach under more realistic detector conditions and high-pileup environments.
Acknowledgements.
We thank Dr Paul Gessinger for helpful contributions on ACTS and Fatras. R. RdA is supported by PID2020-113644GB-I00 from the Spanish Ministerio de Ciencia e Innovación and by PROMETEO/2022/69 from the Spanish GVA.